# Contact non-squeezing in lens spaces

### v1

@@ 2.1. Conventions
2.2. Givental’s non-linear Maslov index
@@
Firstly, let us denote by $(\overline{\phi_{t}})\subset\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\xi_{1})$ the lifted $\mathbb{Z}_{k}$-equivariant contact isotopy starting at the identity associated to the contact isotopy $(\phi_{t})$, i.e. $\pi_{1}^{k}\circ\overline{\phi_{t}}=\phi_{t}\circ\pi_{1}^{k}$ for all $t\in[0,1]$ and $\overline{\phi_{0}}=\mathrm{id}$. We want to construct a path of $C^{1}$-functions $(f_{t}:L_{k}^{2(n+N)+1}\to\mathbb{R})$, for some integer $N\geq 0$, so that for all $t\in[0,1]$
@@
where $g_{t}:\mathbb{S}^{2n+1}\to\mathbb{R}$ is the $\alpha_{1}$-conformal factor of $\overline{\phi}_{t}$, i.e. $\overline{\phi}_{t}^{*}\alpha_{1}=e^{g_{t}}\alpha_{1}$. Extending $\psi_{t}$ to the whole $\mathbb{R}^{2n+2}$ by putting $\psi_{t}(0)=0$ the authors in [21, Prop 2.2 and Prop 2.14] show the existence of an integer $m\in\mathbb{N}$ and a path of $C^{1}$-functions $(F_{t}:\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}\to\mathbb{R})$, where $2N=(2n+2)\cdot 2m$, such that denoting by $\Sigma_{F_{t}}:=\{(x,v)\in\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}\ |\ \partial_{v}F(x,v)\equiv 0\in T^{*}\mathbb{R}^{2n}\}$ the fiber critical points of $F_{t}$ we have444 in the following $\mathbb{R}^{2j}$ is identified with $\mathbb{C}^{j}$, for all $j\in\mathbb{N}_{>0}$, as in the introduction and $T^{*}\mathbb{R}^{2n+2}$ with $T\mathbb{R}^{2n+2}\simeq\mathbb{R}^{2n+2}\times\mathbb{R}^{2n+2}$ using the standard Euclidean metric
@@
Secondly, suppose that $k=p\in\mathbb{N}_{>0}$ is a prime number. If $(\phi^{\prime}_{t})$ is another contact isotopy in $\mathrm{Cont}_{0}(L_{p}^{2n+1},\xi_{p})$ representing the same element $\widetilde{\varphi}$ and if $(f_{t}:L_{p}^{2N+1}\to\mathbb{R})$ and $(f_{t}^{\prime}:L_{p}^{2N^{\prime}+1}\to\mathbb{R})$ are two paths of functions that are induced by two based families of generating functions associated respectively to $(\phi_{t})$ and $(\phi^{\prime}_{t})$, then it is proven in [21, Prop. 2.20 and Section 4] that
@@
2.3. The spectral selector on lens spaces
@@
Remark 2.4.
Lemma 6.2 below states that the conjugation invariance property of the map $\widetilde{\varphi}\mapsto\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$ is actually a consequence of the $\frac{2\pi}{k}$-periodicity of the $\alpha_{k}$-Reeb flow and the $\alpha_{k}$-spectrality of $c_{k}$. This was already noticed by Sandon [32] in the case of the prequantization of the standard symplectic Euclidean space.
@@
For any $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{S}^{1}}(L_{k}^{2n+1},\xi_{k})$, $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
  1. (1)
@@
  2. (2)
    $c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}^{-1})=c_{k}(\widetilde{\varphi})$.
Remark 2.6.
For the purpose of this paper the second statement of Proposition 2.5 is not relevant.
@@
2.5
  1. (1)
    Let us first suppose that $\Pi(\widetilde{\varphi})$ is $\alpha_{k}$-nondegenerate and denote by $T:=c_{k}(\widetilde{\varphi})$, $T^{\prime}:=c_{k}(\widetilde{\psi})$. Then by Lemma 2.7 we have that $\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})=0$ is even. Therefore, since $\widetilde{\psi}$ commutes with $\widetilde{\phi_{\alpha_{k}}^{s}}$ for all $s\in\mathbb{R}$ we have
@@
    The general case can now be deduced by the $C^{1}$-continuity of $c_{k}$ and the $C^{1}$-density, inside $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$, of the set of elements $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ such that $\Pi(\widetilde{\psi})$ is $\alpha_{k}$-nondegenerate (it is a direct consequence of [4, Lemma 3.1] and the fact that $\Pi:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is a local homeomorphism when endowing both spaces with the $C^{1}$-topology).
  2. (2)
    Let $(\psi_{t})$ be a contact isotopy representing $\widetilde{\psi}$ such that $\phi_{\alpha_{k}}^{s}\circ\psi_{t}=\psi_{t}\circ\phi_{\alpha_{k}}^{s}$ for all $t\in[0,1]$ and $s\in\mathbb{R}$. This implies that $\psi_{T}^{*}\alpha_{k}=\alpha_{k}$, for all $T\in[0,1]$, and that
    $\mathrm{Spec}^{\alpha_{k}}(\widetilde{\psi}_{T}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{T}^{-1})=\mathrm{Spec}^{\alpha_{k}}(\widetilde{\varphi})$
    where $\widetilde{\psi}_{T}:=[(\psi_{tT})]$. Moreover it is proved in [4, Lemma 3.2] that $\mathrm{Spec}^{\alpha_{k}}(\widetilde{\varphi})$ is nowhere dense which implies that the map $[0,1]\to\mathbb{R},\ T\mapsto c_{k}(\widetilde{\psi}_{T}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{T}^{-1})\in\mathrm{Spec}^{\alpha_{k}}(\widetilde{\varphi})$ is constant, because continuous, and thus
    $c_{k}(\widetilde{\varphi})=c_{k}(\widetilde{\psi}_{0}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{0}^{-1})=c_{k}(\widetilde{\psi}_{1}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{1}^{-1})=c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}^{-1}).$
@@
A $\mathbb{S}^{1}$-equivariant contactomorphism $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is always $\alpha_{k}$-degenerate since its $\alpha_{k}$-conformal factor is constantly equal to 00 and $(\phi_{\alpha_{k}}^{-T}\circ\phi)^{*}R_{\alpha_{k}}=R_{\alpha_{k}}$ for all $T\in\mathbb{R}$.
3. Contact capacity
@@
The monotonicity property of $C_{k}$ in Theorem 3.1 follows by definition. The invariance by conjugation of $C_{k}$ follows from the invariance by conjugation of the map $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{R},\ \widetilde{\varphi}\mapsto\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$ (see Proposition 2.3). To give the complete proof of Theorem 3.1 it thus remains to prove the last two estimates. Let us admit them for the moment and deduce the proof of Theorem 1.1:
@@
We give the proof of the last two estimates of Theorem 3.1 in Section 5. But before, in the next Section 4 we prove a energy-capacity inequality satisfied by $C_{k}$ which will be crucial to prove the lower bound stated in Theorem 3.1.
4. energy-capacity Inequality
@@
    If $(M,\xi)$ is a closed cooriented contact manifold, the Reeb flow of a generic contact form supporting $\xi$ has a dense orbit [9, Theorem C.4]. Hence, for such contact form $\alpha$, a diffeomorphism $\phi$ cannot displace the closure of the $\alpha$-orbit of any non-empty open set $U$, i.e. $\phi(\overline{\mathcal{O}_{\alpha}(U)})\cap\overline{\mathcal{O}_{\alpha}(U)}\neq\emptyset$.
@@
Let $U\subset L_{k}^{2n+1}$ and $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$. If $\Pi(\widetilde{\psi})$ $\alpha_{k}$-displaces $U$ then $\lceil C_{k}(U)\rceil_{\frac{2\pi}{k}}\leq c_{k}(\widetilde{\psi}^{-1})+\lceil c_{k}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}}$. If moreover $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{S}^{1}}(L_{k}^{2n+1},\xi_{k})$ then $C_{k}(U)\leq c_{k}(\widetilde{\psi}^{-1})+c_{k}(\widetilde{\psi}).$
@@
  1. (1)
    The terminoly "energy-capacity inequality" is borrowed from symplectic geometry where a similar inequality was first described by Hofer and Zehnder while working with the Hofer-Zehnder capacity and the Hofer metric in the standard symplectic Euclidean space [25, Section 5] (see also [36, Prop. 4.12]). The proof of Theorem 4.3 is based on an adaptation of their proof.
  2. (2)
@@
Let $U\subset M$ be a non-empty open set of a closed contact manifold $(M,\ker\alpha)$ and $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}^{c}(U,\ker\alpha)$. If $c:\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)\to\mathbb{R}$ is an $\alpha$-spectral selector and if $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)$ is such that $\Pi(\widetilde{\psi})$ $\alpha$-displaces $U\subset M$ then
@@
Proposition 4.5 is a consequence of the following:
@@
Proof of Lemma 4.6.
Let $t\in\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})$ and $x\in M$ so that $\phi_{\alpha}^{-t}\circ(\psi\circ\phi)$ where $\psi:=\Pi(\widetilde{\psi})$ and $\phi:=\Pi(\widetilde{\varphi})$. In particular $\psi(\phi(x))=\phi_{\alpha}^{t}(x)$. Suppose by contradiction that $x$ lies in the support of $\phi$. Then $\phi(x)\in\mathrm{Supp}(\phi)\subset\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))$ and $\phi_{\alpha}^{t}(x)\in\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))$. But this contradicts the fact that $\psi(\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))\cap\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))=\emptyset$. Therefore $x$ is not in the support of $\phi$. Thus $x$ is a discriminant point of $\phi_{\alpha}^{-t}\circ\psi$ and $t\in\mathrm{Spec}^{\alpha}(\widetilde{\psi})$. In particular $\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})\subset\mathrm{Spec}^{\alpha}(\widetilde{\psi})$. ∎
@@
where the last equality follows from Proposition 4.5. This proves the first statement. Moreover under the assumption that $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{S}^{1}}(L_{k}^{2n+1},\xi_{k})$ one can use Proposition 2.5 to show that the relations in (2) remain true without taking the $\frac{2\pi}{k}$-ceiling function which proves the second statement. ∎

### v2

@@ 2.1. Conventions and definition
2.2. Orderability, Periodic Reeb flow and invariance of spectral selectors
Let $(M,\xi)$ be a closed coorientable contact manifold. From now on we fix a coorientation of $\xi$, i.e. a vector field $X\in\chi(M)$ so that $X(M)\subset TM\setminus\xi$. We say that a contact form $\alpha$ supports $\xi$ if $\ker\alpha=\xi$ and $\alpha(X)>0$.
A contact isotopy $(\phi_{t})$ is said to be non-negative (resp. positive) if for some, and thus all, contact form $\alpha$ supporting $\xi$ we have $\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\geq 0$ (resp. $>0$) for all $x\in M$. Following [16] we define a binary relation on $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ by saying that $\widetilde{\varphi}\preceq\widetilde{\psi}$ if there exists a path $(\widetilde{\varphi}_{t})$ joining them such that $(\Pi(\widetilde{\varphi}_{t}))$ is a non-negative contact isotopy. It is shown in [16] that this binary relation is a partial order if and only there does not exist any positive contractible loop of contactomorphisms, and in this case we say that $(M,\xi)$ is orderable. The following Lemma will be crucial in this paper and was, to some extent, already noticed by Sandon in [36]
Lemma 2.2.
Proof.
3. Spectral selectors for lens spaces
@@
3.0.1. Givental’s non-linear Maslov index
@@
Firstly, let us denote by $(\overline{\phi_{t}})\subset\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\xi_{1})$ the lifted $\mathbb{Z}_{k}$-equivariant contact isotopy starting at the identity associated to the contact isotopy $(\phi_{t})$, i.e. $\pi_{1}^{k}\circ\overline{\phi_{t}}=\phi_{t}\circ\pi_{1}^{k}$ for all $t\in[0,1]$ and $\overline{\phi_{0}}=\mathrm{id}$ where $\pi_{1}^{k}:\mathbb{S}^{2n+1}\to L_{k}^{2n+1}$ is the covering map. We want to construct a path of $C^{1}$-functions $(f_{t}:L_{k}^{2(n+N)+1}\to\mathbb{R})$, for some integer $N\geq 0$, so that for all $t\in[0,1]$
@@
where $g_{t}:\mathbb{S}^{2n+1}\to\mathbb{R}$ is the $\alpha_{1}$-conformal factor of $\overline{\phi}_{t}$, i.e. $\overline{\phi}_{t}^{*}\alpha_{1}=e^{g_{t}}\alpha_{1}$. Extending $\psi_{t}$ to the whole $\mathbb{R}^{2n+2}$ by putting $\psi_{t}(0)=0$ the authors in [22, Prop 2.2 and Prop 2.14] show the existence of an integer $m\in\mathbb{N}$ and a path of $C^{1}$-functions $(F_{t}:\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}\to\mathbb{R})$, where $2N=(2n+2)\cdot 2m$, such that denoting by $\Sigma_{F_{t}}:=\{(x,v)\in\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}\ |\ \partial_{v}F(x,v)\equiv 0\in T^{*}\mathbb{R}^{2n}\}$ the fiber critical points of $F_{t}$ we have666 in the following $\mathbb{R}^{2j}$ is identified with $\mathbb{C}^{j}$, for all $j\in\mathbb{N}_{>0}$, as in the introduction and $T^{*}\mathbb{R}^{2n+2}$ with $T\mathbb{R}^{2n+2}\simeq\mathbb{R}^{2n+2}\times\mathbb{R}^{2n+2}$ using the standard Euclidean metric
@@
Secondly, suppose that $k=p\in\mathbb{N}_{>0}$ is a prime number. If $(\phi^{\prime}_{t})$ is another contact isotopy in $\mathrm{Cont}_{0}(L_{p}^{2n+1},\xi_{p})$ representing the same element $\widetilde{\varphi}$ and if $(f_{t}:L_{p}^{2N+1}\to\mathbb{R})$ and $(f_{t}^{\prime}:L_{p}^{2N^{\prime}+1}\to\mathbb{R})$ are two paths of functions that are induced by two based families of generating functions associated respectively to $(\phi_{t})$ and $(\phi^{\prime}_{t})$, then it is proven in [22, Prop. 2.20 and Section 4] that
@@
3.0.2. The spectral selector on $(L_{k}^{2n+1},\xi_{k})$
@@
For any $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(L_{k}^{2n+1},\xi_{k})$, $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
@@
3.3 Let us first suppose that $\Pi(\widetilde{\varphi})$ is $\alpha_{k}$-nondegenerate and denote by $T:=c_{k}(\widetilde{\varphi})$, $T^{\prime}:=c_{k}(\widetilde{\psi})$. Then by Lemma 3.4 we have that $\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})=0$ is even. Therefore, since $\widetilde{\psi}$ commutes with $\widetilde{\phi_{\alpha_{k}}^{s}}$ for all $s\in\mathbb{R}$ we have
@@
The general case can now be deduced by the $C^{1}$-continuity of $c_{k}$ and the $C^{1}$-density, inside $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$, of the set of elements $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ such that $\Pi(\widetilde{\psi})$ is $\alpha_{k}$-nondegenerate (it is a direct consequence of [4, Lemma 3.1] and the fact that $\Pi:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is a local homeomorphism when endowing both spaces with the $C^{1}$-topology).∎
@@
A $\mathbb{T}^{1}$-equivariant contactomorphism $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is always $\alpha_{k}$-degenerate since its $\alpha_{k}$-conformal factor is constantly equal to 00, i.e. $\phi^{*}\alpha_{k}=\alpha_{k}$, and $(\phi_{\alpha_{k}}^{-T}\circ\phi)^{*}R_{\alpha_{k}}=R_{\alpha_{k}}$ for all $T\in\mathbb{R}$.
4. Contact capacity in Lens spaces
@@
The monotonicity property of $C_{k}$ in Theorem 4.1 follows by definition. The invariance by conjugation of $C_{k}$ follows from the invariance by conjugation of the map $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{R},\ \widetilde{\varphi}\mapsto\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$ (see Lemma 2.2). To give the complete proof of Theorem 4.1 it thus remains to prove the last two estimates. Let us admit them for the moment and deduce the proof of Theorem 1.1:
@@
We give the proof of the last two estimates of Theorem 4.1 in Section 6. But before, in the next Section 5 we prove a capacity-energy inequality satisfied by $C_{k}$ which will be crucial to prove the lower bound stated in Theorem 4.1.
5. Capacity-energy Inequality
@@
    If $(M,\xi)$ is a closed cooriented contact manifold, the Reeb flow of a generic contact form supporting $\xi$ has a dense orbit [10, Theorem C.4]. Hence, for such contact form $\alpha$, a diffeomorphism $\phi$ cannot displace the closure of the $\alpha$-orbit of any non-empty open set $U$, i.e. $\phi(\overline{\mathcal{O}_{\alpha}(U)})\cap\overline{\mathcal{O}_{\alpha}(U)}\neq\emptyset$.
@@
Let $U\subset L_{k}^{2n+1}$ and $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$. If $\Pi(\widetilde{\psi})$ $\alpha_{k}$-displaces $U$ then $\lceil c_{k}(U)\rceil_{\frac{2\pi}{k}}\leq c_{k}(\widetilde{\psi}^{-1})+\lceil c_{k}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}}$. If moreover $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(L_{k}^{2n+1},\xi_{k})$ then $c_{k}(U)\leq c_{k}(\widetilde{\psi}^{-1})+c_{k}(\widetilde{\psi}).$
@@
Let $U\subset M$ be a non-empty open set of a closed contact manifold $(M,\ker\alpha)$ and $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}^{c}(U,\xi)$. If $c:\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)\to\mathbb{R}$ is an $\alpha$-spectral selector and if $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ is such that $\Pi(\widetilde{\psi})$ $\alpha$-displaces $U\subset M$ then
@@
Proposition is a consequence of the following:
@@
Proof.
Let $t\in\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})$ and $x\in M$ a discriminant point of $\phi_{\alpha}^{-t}\circ(\psi\circ\phi)$ where $\psi:=\Pi(\widetilde{\psi})$ and $\phi:=\Pi(\widetilde{\varphi})$. In particular $\psi(\phi(x))=\phi_{\alpha}^{t}(x)$. Suppose by contradiction that $x$ lies in the support of $\phi$. Then $\phi(x)\in\mathrm{Supp}(\phi)\subset\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))$ and $\phi_{\alpha}^{t}(x)\in\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))$. But this contradicts the fact that $\psi(\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))\cap\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))=\emptyset$. Therefore $x$ is not in the support of $\phi$. Thus $x$ is a discriminant point of $\phi_{\alpha}^{-t}\circ\psi$ and $t\in\mathrm{Spec}^{\alpha}(\widetilde{\psi})$. In particular $\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})\subset\mathrm{Spec}^{\alpha}(\widetilde{\psi})$. ∎
@@
where the last equality follows from Proposition 5.5. This proves the first statement. Moreover under the assumption that $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(L_{k}^{2n+1},\xi_{k})$ one can use Proposition 3.3 to show that the relations in (2) remain true without taking the $\frac{2\pi}{k}$-ceiling function which proves the second statement. ∎